# VEGGIE [Uncaptioned image]: Instructional Editing and Reasoning of Video Concepts with Grounded Generation

### v2

@@ Abstract
Recent video diffusion models have enhanced video editing, but it remains challenging to handle instructional editing and diverse tasks (e.g., adding, removing, changing) within a unified framework. In this paper, we introduce VEGGIE, a Video Editor with Grounded Generation from Instructions, a simple end-to-end framework that unifies video concept editing, grounding, and reasoning based on diverse user instructions. Specifically, given a video and text query, VEGGIE first utilizes an MLLM to interpret user intentions in instructions and ground them to the video contexts, generating frame-specific grounded task queries for pixel-space responses. A diffusion model then renders these plans and generates edited videos that align with user intent. To support diverse tasks and complex instructions, we employ a curriculum learning strategy: first aligning the MLLM and video diffusion model with large-scale instructional image editing data, followed by end-to-end fine-tuning on high-quality multitask video data. Additionally, we introduce a novel data synthesis pipeline to generate paired instructional video editing data for model training. It transforms static image data into diverse, high-quality video editing samples by leveraging Image-to-Video models to inject dynamics. VEGGIE shows strong performance in instructional video editing with different editing skills, outperforming the best instructional baseline as a versatile model, while other models struggle with multi-tasking. VEGGIE also excels in video object grounding and reasoning segmentation, where other baselines fail. We further reveal how the multiple tasks help each other and highlight promising applications like zero-shot multimodal instructional and in-context video editing.
@@ 1 Introduction
To address these challenges, we introduce VEGGIE, a Video Editor with Grounded Generation from Instructions. VEGGIE unifies video concept grounding and editing without relying on additional layout, mask guidance, or intermediate caption [77, 74, 35, 41, 78]. Instead, we formulate the problem as end-to-end grounded generation in pixel space, using only a diffusion loss. Specifically, given a video and a text query, VEGGIE first leverages an MLLM to interpret complex instructions, generating frame-wise conditions. Unlike prior methods [74, 42] that use discrete text tokens as conditions, which disconnect the pipeline and block gradient propagation, VEGGIE employs continuous, learnable task query embeddings per frame. This enables end-to-end training and effectively captures grounded task representations for diffusion model conditioning. To handle diverse tasks and accurately interpret complex queries, we employ a curriculum learning strategy that begins by aligning MLLMs with diffusion models using massive paired instructional image editing data and then fine-tuning the model end-to-end on high-quality multitask video data to adapt video. Unlike tool-use methods [74, 16, 35], VEGGIE formulates both video grounding and instructional editing in the same video-to-video task formulation, enabling efficient handling through a unified single model. To further support end-to-end training, we introduce a novel automatic instructional video data generation pipeline that lifts high-quality instructional image editing data into the video domain using image-to-video and video evaluation tools.
Existing video editing benchmarks do not provide wide and uniform coverage of diverse editing skills [67, 58]. To address this gap, we contribute VEG-Bench, an instructional video editing benchmark that spans 8 editing skills: concept addition, removal, object changing, environment background changing, visual feature changing, stylization, object grounding, and reasoning segmentation. Each skill is evaluated using a dedicated suite of metrics. We assess the proposed VEGGIE alongside 6 baselines on VEG-Bench. VEGGIE demonstrates strong performance across diverse editing skills, outperforming the best instructional baseline as a versatile, all-in-one model, while other models struggle with multi-tasking. Additionally, VEGGIE excels in video object grounding and reasoning segmentation tasks, where other baselines fall short. We show further analysis of how multi-task learning enhances our framework and highlight applications such as zero-shot multimodal instructional following and few-shot in-context editing. Our contributions are summarized as follows:
@@ 3 Our Method: VEGGIE
Figure 1: Overview of our proposed end-to-end VEGGIE framework. Our Multimodal Large Language Model first understands input video frames and diverse user instructions, then it generates frame-wise reasoning queries that maintain per-frame editing conditions for the video diffusion model. The video diffusion model will render the MLLM-generated conditions to the pixel space for diverse tasks, including video editing (add/remove/change), video grounding, and video reasoning segmentation with questions. We only apply diffusion loss for the whole pipeline training.
In this paper, we introduce a Video Editor with Grounded Generation from Instructions (VEGGIE), a unified and versatile generative video model. It combines the complex instruction understanding and reasoning capabilities of MLLMs with the generative capacity of VidDMs. The model is trained end-to-end with diffusion loss only. VEGGIE efficiently handles diverse user inputs, including direct instructions, complex questions requiring in-depth reasoning, and multimodal conditioning. It performs various pixel-level manipulations, enabling tasks such as video concept addition, removal, changing, stylization, grounding, and reasoning segmentation based on user instructions. We elaborate on the model design (Sec. 3.1), training and inference process (Sec. 3.2), and data curation (Sec. 3.3),.
@@ 3.1 Model Architecture
VEGGIE consists of four main components (see Fig. 1): (1) a multimodal large language model, (2) a set of learnable grounded task queries, (3) an alignment network (single-layer MLP) that projects the MLLM output into the condition space of the diffusion model, and (4) a video diffusion model initialized from an instructional image editing model [82]. Our model first generates latent conditions for target video frames by querying multimodal context using an MLLM, then renders these conditions at the pixel level through a video diffusion model, as detailed below.
MLLM for Generating Grounded Task Guidance. As illustrated in the left of Fig. 1, given a video consisting of a sequence of frames $V=[f_{1},\dots,f_{n}]$, where $n$ is the frame number of the given video, a user instruction/question $I$, our goal is to obtain the response $\widehat{V}=[\widehat{f_{1}},\dots,\widehat{f_{n}}]$ at pixel space that faithfully reflects user instruction about the given video. The MLLM module processes both the input video $V$ and a user instruction $I$ to generate a sequence of grounded task tokens per frame: $C=[c_{1},\dots,c_{n}]$, which are input and output in parallel. These tokens serve as task guidance and implicitly encode the target manipulation, such as object attributes, spatial relationships, or style transfer parameters. The MLLM ensures the model captures both explicit user instructions and implicit reasoning needs.
VidDM for Rendering MLLM Guidance at Pixel Space. As illustrated in the right of Table 1, the VidDM takes the original video $V$ and the grounded task tokens $C$ as conditions to synthesize the target video $\widehat{V}$. The original video is concatenated with the noise volume, and the task tokens are input to the cross-attention. With grounded task guidance in denoising steps, the generation process ensures that the output faithfully follows user instructions while preserving the video’s structure and motion dynamics. Through iterative denoising, it refines each frame while maintaining temporal consistency, applying pixel modifications coherently for a smooth and visually consistent output video $\widehat{V}$.
@@ 3.2 Curriculum Learning from Images to Videos
where $\theta$ represents the model parameters, $C_{T}$ and $C_{V}$ denote the task tokens and video conditions, $\varnothing$ is the null value, $z_{t}$ is the noised latent at timestamp $t$, and $g_{T}$ and $g_{V}$ are the task guidance and video guidance scales, respectively. More training details are included later in Sec. 4.2.
@@
Table 1: Summary of our data for VEGGIE training. R.: Reasoning, E.: Editing, G.: Grounding. # Img/Vid denotes the number of images or videos, and # Ins. represents the number of instruction-image/video pairs.
@@ 3.3 Data Curation Pipeline
Existing video editing models, both instructional and non-instructional, struggle with diverse editing skills due to the lack of high-quality multitasking fine-tuning data. In this section, we introduce our data curation strategy to support VEGGIE in achieving versatile video editing skills. As listed in Tab. 1, we collect 3.4M image and 133.9K video data from diverse sources to support our VEGGIE curriculum learning as discussed in Sec. 3.2. We create our training dataset from two sources: (1) collecting existing image and video data and converting it into an instructional editing format, and (2) synthesizing new instructional video editing samples using existing datasets and generative models.
@@
Synthesizing Instructional Video Editing Data via Image-to-Video Animation. Recent methods [53, 3] generate synthetic instructional video-editing data by first creating text instructions with LLM, then getting edited videos via T2V models and prompt-to-prompt editing [23]. While these methods adapt image-based editing pipelines [9] to videos, the generated data suffer from temporal-consistency issues. To address this gap, we propose a novel image-to-video animation strategy that leverages the abundance of high-quality image-level instructional editing datasets [82, 63], which provide well-annotated instructions, paired edited images, and well-organized editing skill categories. As illustrated in Fig. 2, given an original image $I$, an edited image $\bar{I}$, and an instruction from an instructional image editing dataset [63], our approach involves three key steps. First, we use an offline MLLM [71, 60] to generate an image caption and an animation prompt that describes plausible motion within the image. Next, an image-to-video (I2V) model animates the image into a video $V$. Finally, we generate the corresponding edited video $\bar{V}$ using a first-frame-conditioned video editing model [35], leveraging $\bar{I}$ as a strong prior to ensure consistent edits across frames. Finally, to ensure data quality, we evaluate each original-edited video pair with automatic video quality evaluation metrics [30], which assess the generated videos from diverse dimensions, e.g., motion smoothness, image quality, and background consistency. This pipeline transforms carefully curated image-based datasets into instructional video-editing resources while preserving the precision of the original edits. As a result, our data method expands the availability of high-quality synthetic video-editing data, supporting a wider range of editing tasks in our end-to-end unified framework. More details on data generation, prompting, examples, and pre/post-processing, are in Sec. 4.2 and the Appendix.
@@ 4 Experiments
Figure 3: Qualitative comparison of editing results across 8 different abilities (splitting visual features into color and texture). We provide zoom-in details for a more detailed comparison. Best viewed in color. More in Appendix. We provide videos in the Supplementary Materials.
@@
Table 2: Comparison of video editing task with instructional / non-instructional models on VEG-Bench. --: the task is not capable of non-instructional models. We gray out numbers of non-instructional models that are in different categories.
@@
We first introduce the VEG-Bench Benchmark and then demonstrate the superiority of VEGGIE across diverse video instructional editing skills. More visualization, evaluation, and implementation details are in the Appendix.

### v3

# VEGGIE [Uncaptioned image]: Instructional Editing and Reasoning Video Concepts with Grounded Generation

@@ Abstract
While recent video diffusion models enable video editing, unifying diverse instructional editing tasks (e.g., add, remove, modify) under a single framework remains a significant challenge. In this paper, we introduce VEGGIE, a Video Editor with Grounded Generation from Instructions, a simple end-to-end framework that unifies video concept editing, grounding, and reasoning based on diverse user instructions. Specifically, given a video and text query, VEGGIE first utilizes an MLLM to interpret user intentions in instructions and ground them to the video contexts, generating frame-specific grounded task queries for pixel-space responses. A diffusion model then renders these plans and generates edited videos that align with user intent. To support diverse tasks and complex instructions, we employ a curriculum learning strategy: first aligning the MLLM and video diffusion model with large-scale instructional image editing data, followed by end-to-end fine-tuning on high-quality multitask video data. Additionally, we introduce a novel data synthesis pipeline to generate paired instructional video editing data for model training. It transforms static image data into diverse, high-quality video editing samples by leveraging Image-to-Video models to inject dynamics. VEGGIE shows strong performance in instructional video editing with different editing skills, outperforming the best instructional baseline as a versatile model, while other models struggle with multi-tasking. VEGGIE also excels in video object grounding and reasoning segmentation, where other baselines fail. We further reveal how the multiple tasks help each other and highlight promising applications like zero-shot multimodal instructional and in-context video editing.
@@ 1 Introduction
To address these challenges, we introduce VEGGIE, a Video Editor with Grounded Generation from Instructions. VEGGIE unifies video concept grounding and editing without relying on additional layout, mask guidance, or intermediate caption [78, 75, 35, 42, 79, 39]. Instead, we formulate the problem as end-to-end grounded generation in pixel space, using only a diffusion loss. Specifically, given a video and a text query, VEGGIE first leverages an MLLM to interpret complex instructions, generating frame-wise conditions. Unlike prior methods [75, 43] that use discrete text tokens as conditions, which disconnect the pipeline and block gradient propagation, VEGGIE employs continuous, learnable task query embeddings per frame. This enables end-to-end training and effectively captures grounded task representations for diffusion model conditioning. To handle diverse tasks and accurately interpret complex queries, we employ a curriculum learning strategy that begins by aligning MLLMs with diffusion models using massive paired instructional image editing data and then fine-tuning the model end-to-end on high-quality multitask video data to adapt video. Unlike tool-use methods [75, 16, 35], VEGGIE formulates both video grounding and instructional editing in the same video-to-video task formulation, enabling efficient handling through a unified single model. To further support end-to-end training, we introduce a novel automatic instructional video data generation pipeline that lifts high-quality instructional image editing data into the video domain using image-to-video and video evaluation tools.
Existing video editing benchmarks do not provide wide and uniform coverage of diverse editing skills [68, 59]. To address this gap, we contribute VEG-Bench, an instructional video editing benchmark that spans 8 editing skills: concept addition, removal, object changing, environment background changing, visual feature changing, stylization, object grounding, and reasoning segmentation. Each skill is evaluated using a dedicated suite of metrics. We assess the proposed VEGGIE alongside 6 baselines on VEG-Bench. VEGGIE demonstrates strong performance across diverse editing skills, outperforming the best instructional baseline as a versatile, all-in-one model, while other models struggle with multi-tasking. Additionally, VEGGIE excels in video object grounding and reasoning segmentation tasks, where other baselines fall short. We show further analysis of how multi-task learning enhances our framework and highlight applications such as zero-shot multimodal instructional following (Fig.˜1) and few-shot in-context editing (Fig.˜2). Our contributions are summarized as follows:
@@ 3 Our Method: VEGGIE
Figure 3: Overview of our proposed end-to-end VEGGIE framework. Our Multimodal Large Language Model first understands input video frames and diverse user instructions, then it generates frame-wise reasoning queries that maintain per-frame editing conditions for the video diffusion model. The video diffusion model will render the MLLM-generated conditions to the pixel space for diverse tasks, including video editing, video grounding, and video reasoning segmentation with questions. We only apply diffusion loss for the whole pipeline training.
In this paper, we introduce a Video Editor with Grounded Generation from Instructions (VEGGIE), a unified and versatile generative video model. It combines the complex instruction understanding and reasoning capabilities of MLLMs with the generative capacity of VidDMs. The model is trained end-to-end with diffusion loss only. VEGGIE efficiently handles diverse user inputs, including direct instructions, complex questions requiring in-depth reasoning, and multimodal conditioning. It performs various pixel-level manipulations, enabling tasks such as video concept addition, removal, changing, stylization, grounding, and reasoning segmentation based on user instructions. We elaborate on the model design (Sec.˜3.1), training and inference process (Sec.˜3.2), and data curation (Sec.˜3.3),.
@@ 3.1 Model Architecture
VEGGIE consists of four main components (see Fig.˜3): (1) a multimodal large language model, (2) a set of learnable grounded task queries, (3) an alignment network (single-layer MLP) that projects the MLLM output into the condition space of the diffusion model, and (4) a video diffusion model initialized from an instructional image editing model [83]. Our model first generates latent conditions for target video frames by querying multimodal context using an MLLM, then renders these conditions at the pixel level through a video diffusion model, as detailed below.
MLLM for Generating Grounded Task Guidance. As illustrated in the left of Fig.˜3, given a video consisting of a sequence of frames $V=[f_{1},\dots,f_{n}]$, where $n$ is the frame number of the given video, a user instruction/question $I$, our goal is to obtain the response $\widehat{V}=[\widehat{f_{1}},\dots,\widehat{f_{n}}$] at pixel space that faithfully reflects user instruction about the given video. The MLLM module processes both the input video $V$ and a user instruction $I$ to generate a sequence of grounded task tokens per frame: $C=[c_{1},\dots,c_{n}]$, which are input and output in parallel. These tokens serve as task guidance and implicitly encode the target manipulation, such as object attributes, spatial relationships, or style transfer parameters. The MLLM ensures the model captures both explicit user instructions and implicit reasoning needs.
VidDM for Rendering MLLM Guidance at Pixel Space. As illustrated in the right of Table˜3, the VidDM takes the original video $V$ and the grounded task tokens $C$ as conditions to synthesize the target video $\widehat{V}$. The original video is concatenated with the noise volume, and the task tokens are input to the cross-attention. With grounded task guidance in denoising steps, the generation process ensures that the output faithfully follows user instructions while preserving the video’s structure and motion dynamics. Through iterative denoising, it refines each frame while maintaining temporal consistency, applying pixel modifications coherently for a smooth and visually consistent output video $\widehat{V}$.
@@ 3.2 Curriculum Learning from Images to Videos
where $\theta$ represents the model parameters, $C_{T}$ and $C_{V}$ denote the task tokens and video conditions, $\varnothing$ is the null value, $z_{t}$ is the noised latent at timestamp $t$, and $g_{T}$ and $g_{V}$ are the task guidance and video guidance scales, respectively. More training details are included later in Appendix.
@@
Table 1: Summary of our data for training. R.: Reasoning, E.: Editing, G.: Grounding. # Img/Vid: the number of images/videos, and # Ins.: the number of instruction-image/video pairs.
@@ 3.3 Data Curation Pipeline
Existing video editing models, both instructional and non-instructional, struggle with diverse editing skills due to the lack of high-quality multitasking fine-tuning data. In this section, we introduce our data curation strategy to support VEGGIE in achieving versatile video editing skills. As listed in Tab.˜1, we collect 3.4M image and 133.9K video data from diverse sources to support our VEGGIE curriculum learning as discussed in Sec.˜3.2. We create our training dataset from two sources: (1) collecting existing image and video data and converting it into an instructional editing format, and (2) synthesizing new instructional video editing samples using existing datasets and generative models.
@@
Synthesizing Instructional Video Editing Data via Image-to-Video Animation. Recent methods [54, 3] generate synthetic instructional video-editing data by first creating text instructions with LLM, then getting edited videos via T2V models and prompt-to-prompt editing [23]. While these methods adapt image-based editing pipelines [9] to videos, the generated data suffer from temporal-consistency issues. To address this gap, we propose a novel image-to-video animation strategy that leverages the abundance of high-quality image-level instructional editing datasets [83, 64], which provide well-annotated instructions, paired edited images, and well-organized editing skill categories. As illustrated in Fig.˜4, given an original image $I$, an edited image $\bar{I}$, and an instruction from an instructional image editing dataset [64], our approach involves three key steps. First, we use an offline MLLM [72, 61] to generate an image caption and an animation prompt that describes plausible motion within the image. Next, an image-to-video (I2V) model animates the image into a video $V$. Finally, we generate the corresponding edited video $\bar{V}$ using a first-frame-conditioned video editing model [35], leveraging $\bar{I}$ as a strong prior to ensure consistent edits across frames. Finally, to ensure data quality, we evaluate each original-edited video pair with automatic video quality evaluation metrics [30], which assess the generated videos from diverse dimensions, e.g., motion smoothness, image quality, and background consistency. This pipeline transforms carefully curated image-based datasets into instructional video-editing resources while preserving the precision of the original edits. As a result, our data method expands the availability of high-quality synthetic video-editing data, supporting a wider range of editing tasks in our end-to-end unified framework. More details on data generation, prompting, examples, and pre/post-processing are in the Appendix.
@@ 4 Experiments
Figure 5: Qualitative comparison of editing results across 8 different abilities (splitting visual features into color and texture). We provide zoom-in details for a more detailed comparison. Best viewed in color. More in Appendix.
@@
Table 2: Comparison of video editing task with instructional / non-instructional models on VEG-Bench. -: the task is not capable of non-instructional models. We gray out numbers of non-instructional models that are in different categories.
@@
We first introduce the VEG-Bench Benchmark and then demonstrate the superiority of VEGGIE across diverse video instructional editing skills. More experiments, visualization, and implementation details are in the Appendix.
4.1 VEG-Bench and Metrics
As no existing benchmark is designed for fine-grained instructional video editing skills, we manually collect and annotate VEG-Bench, containing 132 video-instruction pairs that balanced cover 8 different video generative skills (15-20 for each). Beyond standard metrics, including text-to-video alignment (CLIP-Text [55]), video smoothness (CLIP-F [55]), and image quality (MUSIQ [32]), we also first introduce MLLM-as-a-Judge to give a holistic evaluation score according to the given original video, edited video, and user instruction. It is achieved by prompting GPT-4o [21] to evaluate whether the requested semantic change has been fulfilled, using a scale from 1 to 10. For addition and removal, we also introduce an object detector (GroundingDiNo [45]) to detect if the object is added/removed faithfully. For grounding and reasoning segmentation, we following video grounding tasks [12, 33, 58, 14] and adopt the Jaccard index ($\mathcal{J}$[31], F-measure ($\mathcal{F}$[13], and their mean ($\mathcal{J\&F}$). We also compute SSIM between the generated video and the original video masked with GT masks. More evaluation/metrics details are included in Appendix.
Settings Removal (FVD $\downarrow$) Grounding (SSIM $\uparrow$)
Grd.-only - 52.34
Rmv.-only 1098.52 -
Mixed 987.80 55.21
Figure 6: Comparison between single- and multi-skill models with different data training. We find tasks can help each other.
4.2 Experimental Results
Instructional Video Editing over Diverse Skills. As shown in Tab.˜2, we evaluate 7 different models on VEG-Bench across 8 distinct editing skills. Overall, VEGGIE demonstrates the best performance among instructional video editing models. Compared to VEGGIE, non-instructional models often struggle with concept removal and addition. This limitation arises because these models rely on attention control or additional conditions (e.g., depth maps) that impose strong priors, constraining the model and making object addition or removal challenging. We also observe that InsV2V achieves high scores in quality and smoothness metrics, but underperforms in alignment and MLLM judgment, which demand faithful semantic changes. Qualitative examples in Fig. 5 illustrate that InsV2V often makes minimal changes to the input video, resulting in high video quality but unfaithful outputs. In contrast, VEGGIE strikes a better balance, delivering both high-quality visuals and accurate semantic alignment with the intended edits.
Can Multi-Task Help Each Other? To test the previous hypothesis, we train our model on the VPLM [75] dataset, which includes paired grounding and removal tasks (approximately 5.5K samples for each task). We focus on these tasks as representative examples due to their straightforward evaluation against ground truth. As shown in Table˜3, multitask training yields a lower FVD score and a higher SSIM score, demonstrating that learning to locate and remove a video concept can mutually reinforce performance. We show an example in Fig. 6. However, this conclusion only holds with a balanced data combination. We also observe that an excessive amount of grounding data can introduce more artifacts and negatively impact visual editing skills.
5 Conclusion
We present VEGGIE, a unified end-to-end model for instructional video editing that handles diverse pixel-level tasks. VEGGIE leverages MLLM for robust instruction understanding and employs a video diffusion model to execute pixel-level edits. Our framework uses a single diffusion loss for end-to-end optimization across varied tasks/skills. We also introduce a novel synthetic data generation pipeline and VEG-Bench, a benchmark that assesses a broad range of editing skills. Our VEGGIE outperforms previous methods as a versatile, all-in-one solution. We hope our model, data, and benchmark to advance research on instructional generative video models.
Supplementary Material
6 Appendix
In this Appedix, we provide extra details on
  • Limitation and future work discussion.
6.1 Implementation Details
Training Details. Our MLLM is initialized with LLaVA-OneVision-7B (LLaVA-OV) [38]. Our VidDM is initialised from the instructional image editing model, MagicBrush [83] with Stable Diffusion v1.5 backbone [57]. We further inflated 2D convolution layers with temporal attention layers, following AnimateDiff [22] to adapt videos. Our VEGGIE adopts a 2-stage curriculum training strategy (Sec.˜3.2). In the first stage, we fully fine-tune the 2D convolution layers in the UNet, the alignment network, and the task query tokens in the MLLM on image data, with 862M trainable parameters. In the second stage, we train all 3 dimensions in the UNet, the alignment network, the task query tokens, and a LoRA in the MLLM, leading to 1.3B trainable parameters. Both stages are trained end-to-end with only a diffusion loss. More details are in the Appendix.
We keep the VAE encoder and decoder frozen during the entire training process. In the first stage, we keep the MLLM (including visual encoder, MLP projector, and LLM) frozen, and fully fine-tune learnable grounded task queries, alignment network, and diffusion model, leading to around 800M training parameters. We set $1e^{-4}$ learning rate, and 96 batch size on each GPU. We use 16 A100 GPUs for the first stage of fine-tuning with 25K steps. In the second stage, we insert LoRA [27] modules into the LLM backbone, and inflate diffusion models by inserting extra temporal layers as in AnimateDiff [22]. We fine-tune LoRA, alignment network, learnable grounded task query tokens, and the diffusion model, leading to around 1.3B trainable parameters. We set $5e^{-4}$ learning rate, and 1 batch size with 8 gradient accumulation steps on 32 A100 GPUs. For LoRA, we set lora rank 64, lora alpha 16, and lora dropout 0.05. We train the second stage video model 2.5K step with 8 uniformly sampled frames.
Evaluation and Baseline Details. We primarily compare our model with strong instructional editing models [19, 9, 66]. Additionally, we include non-instructional editing models [41, 10, 20] for completeness, although these are not fair baselines since they are not end-to-end and rely on additional conditions, such as depth maps or intermediate captions.
We randomly sample 3 seeds for both our method and baseline methods. In our experiments, we use different classifier-free guidance scores ($g_{T}$ and $g_{V}$ in Sec.˜3.2) for different skills. Specifically, we set $g_{T}=14.5$ and $g_{V}=1.5$ for grounding and reasoning segmentation, while for other editing skills, we use $g_{T}=10.5$ and $g_{V}=2.0$.
We compare the model judged best for each video sample. The agreement between human and MLLM judgments is 0.74, whereas the agreement between human and CLIP is only 0.45. We conducted 5 times of the MLLM evaluation and took an average.
6.2 Data Collection Details
As mentioned in the earlier Sec.˜3.3, beyond collecting existing data, we proposed a novel data synthesis pipeline to generate instructional video data by animating images in the instructional image dataset.
We first use QWen2-VL [61] to caption the original image and give an animation prompt to animate the image via CogVideX1.5-I2V [74]. Please refer Tab.˜5 and Tab.˜6 to our prompt for caption and animation. After getting the animated video, we utilize AnyV2V [35] to edit the video based on the reference image (edited image from image dataset). The reference image gives a strong prior to maintaining the image dataset’s high-quality edit and thus transfer it to the video via the video editing model.
Next, we filter out videos by evaluating VBench metrics [30], including aesthetic quality, motion smoothness, image quality, subject consistency, and background consistency. We set thresholds at 0.6 for aesthetic quality, 65 for imaging quality, 0.9 for motion smoothness, subject consistency, and background consistency. We provide our generated data visualization in Fig.˜8.
6.3 More Quantative Results & Discussion